# Unlocking Doping Effects on Altermagnetism in MnTe: Emergence of Quasi-altermagnetism

### v2

@@ III.4 Quasi-altermagnetism
Quasi-altermagnetism, as we define, refers to a magnetic phase that emerges when an altermagnet loses its rotational symmetries. This loss of symmetry can occur due to external stimuli such as strain, electric field, defects, and vacancies or chemical doping, as is the case in the present study. Like the altermagnets, quasi-altermagnets lack both translational symmetry and inversion symmetry. They do not exhibit Zeeman splitting as in the ferromagnets. Instead, they show an anisotropic momentum-dependent spin-splitting as in altermagnets. However, the splitting is not equal and opposite on either side of the nodal plane. In fact, the nodal plane itself is blurred and ill-defined in the case of quasi-altermagnets due to the fact that the sub-band degeneracy may get lifted on this plane. Furthermore, if the system is an insulator, the net magnetization remains zero, like the antiferromagnets and altermagnets. However, if it is metallic, it can have uncompensated magnetization. For example, MnTe1-xSex with MSG symmetry $C2^{\prime}/m^{\prime}2$ (No.12.62) and $P\bar{3}m^{\prime}1$ (No.164.89) are insulators which exhibit quasi-altermagnetism with zero net magnetization. However, MnTe1-xSbx and MnTe1-xIx with the same MSG symmetry are quasi-altermagnetic metals, and these compounds have uncompensated magnetization. Fig. 5 illustrates the key features of quasi-altermagnetism and how they differ from ferromagnetism, antiferromagnetism, and altermagnetism.
While, altermagnetism requires perfect rotational symmetries, quasi-altermagnetism is a perturbative deviation effect. The constant energy plots of different pairs of bands in altermagnetic and quasi-altermagnetic materials, shown in Fig. 6, illustrates that quasi-altermagnetism originates from the breakdown in rotational symmetry of altermagnets, rather that the superposition of different types of altermagnetic order in varied proportion. In most of the practical systems, it is therefore highly likely that quasi-altermagnetism exists instead of altermagnetism.
@@ III.5 Anomalous Hall effect in MnTe with paired non-magnetic atom substitutions
The anomalous Hall effect (AHE) is a phenomenon that arises due to broken TRS and strong SOC. Altermagnets having strong SOC can exhibit AHC as the TRS is broken. However, the existence of AHE in altermagnets is not inherently guaranteed since it crucially depends on the orientation of the Néel vector. Hexagonal MnTe is one of the prominent altermagnetic materials in which AHE is both theoretically predicted and experimentally observed [19, 21]. But this is only when the Néel vector lies in the basal plane. When Néel vector is oriented along the out-of-plane direction, pristine MnTe does not exhibit AHE. In this configuration, AHC is symmetry-forbidden due to the presence of operations such as the three-fold rotations $C^{\pm}_{3z}$ and in-plane two-fold rotations $C_{210}$ and $C_{120}$ consistent with MSG $P6_{3}^{\prime}/m^{\prime}m^{\prime}c$ (No. 194.268) [21].
In this section, we investigate the potential for generating a non-zero AHC in MnTe with Néel vector along the out-of-plane direction, through doping induced symmetry breaking. Similar to the pristine case, MnTe with either a single (non-magnetic) substitution or a pair (non-magnetic) substitution configuration with MSG $P\bar{6}^{\prime}m^{\prime}2$ (No.187.211) does not exhibit AHE. It is because the symmetry operations that prohibit AHE remain preserved in these configurations. However, in the other pair doping configurations with reduced symmetry, AHE becomes symmetry allowed, leading to a finite AHE. For the pair doped configuration with MSG $P\bar{3}m^{\prime}1$ (No.164.89), a Hall pseudovector along the $c$-axis becomes symmetry-allowed, as the orthogonal rotation axes combined with TRS ($\tau C_{2x}$, $\tau C_{2y}$) leave it invariant. In lower-symmetry structures with MSGs $Fm^{\prime}m^{\prime}2$ (No.42.222) and $Pm^{\prime}m^{\prime}2$ (No.25.60), the presence of the $C_{2z}$ symmetry operation enforces the vanishing of AHC components $\sigma_{yz}$ and $\sigma_{xz}$, while imposing no constraints on $\sigma_{xy}$. On the other hand, the lowest-symmetry and the most probable configuration with MSG $C2^{\prime}/m^{\prime}2$ (No.12.62) lacks any pure rotational symmetry and retains only the combined symmetry operation $\tau C_{2y}$, which permits non-zero components $\sigma_{yz}$ and $\sigma_{xy}$.
@@
In Fig. 7, we show the non-zero components of AHC and $\Omega$ for candidate materials in two distinct classes of pair non-magnetic atom doped MnTe systems – an altermagnet and a quasi-altermagnet – each incorporating Sb and I atoms as dopants. From the figure, we find that for MSG 25.60 (both in Sb-doped [in (e)] and I-doped [in (g)] cases), the net contribution coming from $\Omega_{x}$ and $\Omega_{y}$ vanishes, as there exists always a negative counterpart for each of the positive Berry curvature. In contrast, $\Omega_{z}$ component has a net contribution over the entire Brillouin zone–contributing to non-zero $\sigma_{xy}$ as obtained from equation 1.
In a similar way, for MSG. 12.62, in Fig. 7 (f) and (i), one can note that, only $\Omega_{y}$ component gets compensated, and other two ($\Omega_{x}$ and $\Omega_{z}$) remain uncompensated throughout the Brillouin zone. This gives rise to two non zero components of AHC – $\sigma_{yz}$ and $\sigma_{xy}$ – as depicted in Fig. 7 (b) and (d), respectively.
@@ III.6 Microscopic origin of altermagnetism in doped MnTe: A model Hamiltonian approach
where, i,ji,\ j are site indices, $\mu,\ \nu$ depict orbital indices, and $\sigma$ describes the spin indices. $\tau=1$ represents the sublattice where spin-up states form the majority spin channel, and $\tau=2$ represents the sublattice where spin-down states form the majority spin channel. The detailed analysis for the chemical bonding origin of altermagnetism in NiAs prototype compounds can be found in Ref. [25]. We primarily focus on the pair doping cases, as single doping is a subset of this. Here, we intend to see the effect of doping in MnTe, which will affect only a part of the existing chemical bonding of the pristine crystal.
@@
If we examine the crystal structure of MnTe (for all NiAs prototypes) it can be presented as a series of chainlike interaction blocks, as shown in Figs. 8(a) and 9(a). In the figures, each coloured cuboid within the unit cell represents one such interaction block. If the electron-electron hopping interactions are restricted to the second nearest neighbour, these blocks are isolated from each other to behave like a ”quasi-1D” chain. Therefore, the up- and down-spin blocks of the total Hamiltonian matrix can be written in terms of a set of block diagonal matrices with each block representing one of this quasi-1D chain. We note that, by ignoring the interaction beyond the second neighbour the altermagnetic characteristics are not affected  [25]. As the sole objective is to demonstrate how altermagnetic band dispersion gets affected in the presence of doping, for simplicity, we adapted a single orbital model (i.e., one $d$-orbital from the magnetic atom and one $p$-orbital from the non-magnetic atom). Earlier we have exclusively shown that it is possible to demonstrate altermagnetism using a single orbital picture. Within this framework, the basis of the Hamiltonian is now formed by the two $d$-orbitals, one each from the opposite spin sublattices, representing the magnetic site, and two effective hybridized orbitals representing the non-magnetic site. There are three nearest neighbour non-magnetic isoplanar atoms forming a triangle, one below a magnetic site and one above it [see right panel Fig. 8 (a)]. For each of the triangles, one effective hybridized orbital is formed out of the linear combination of three identical $p$-orbitals, coming from the non-magnetic atoms.
@@
Figure 8: SK-TB analysis of altermagnetism in MnTe with pair doping. The primitive unit cell of the representative crystal of MSG $P\bar{6}^{\prime}m^{\prime}2$ (No.187.211) [see Fig. 3(a)] is considered here. (a) Identification of different quasi-1D interaction blocks shown by colour cuboids. From the electronic structure point of view, they are isolated from each other, if the hopping interactions are restricted up to the second nearest neighbours. In each quasi-1D block, a magnetic atom is capped by two iso-planar triangles, one above and one below, formed by the non-magnetic atoms. For the single-doped case, there exist four such quasi-1D interaction blocks, out of which one remains unaffected from doping (shown in grey) and the other three contain one doped atom Se/Sb/I. (b) Schematic illustration of the block diagonal tight-binding model Hamiltonian with each block representing one of the quasi-1D chains. (c) A visualization of the hopping interactions in the Hamiltonian matrix of one of the quasi-1D chains. (d) The band structure of configuration (c) is shown, where the altermagnetism is not affected by doping, with the characteristic AMSS retained.
@@
Figure 9: SK-TB analysis of quasi-altermagnetism in MnTe with pair doping. The quasi-1D interaction blocks are designed for the representative crystal of MSG $C2^{\prime}/m^{\prime}$ (No.12.62) as depicted in Fig. 3(b) to study the effect of doping. (a) There exist four different interaction blocks in the primitive unit cell in the framework of up to second neighbour interactions, which are shown by the colour cuboids. Out of them, two interaction blocks contain only one doped atom (as shown by green and magenta colours in the upper panel), and the remaining two interaction blocks contain two doped atoms (as shown by brown and blue colours in the middle panel). While in the former case, the altermagnetic band structure is retained due to the presence of mirror symmetry, it vanishes, and a quasi-altermagnetic band structure is formed in the latter case because of the breaking of mirror symmetry. (b) Schematic illustration of the block diagonal tight-binding model Hamiltonian, with each block representing one of the quasi-1D chains. (c) The various hopping interactions are presented for a pair doped interaction block. Quasi-altermagnetic band structure for this configuration when there is a difference between the nonmagnetic onsite energies (d), hopping strengths (e), and local spin moments (f). In a real system, all three of them contribute together.
@@
Case I: Altermagnetism– Out of five different MSG families of pair doped configurations, $P\bar{6}^{\prime}m^{\prime}2$ (No.187.211), $Fm^{\prime}m^{\prime}2$ (No.42.222), and $Pm^{\prime}m^{\prime}2$ (No.25.60) come under the altermagnetism category, where a primitive unit cell exists with half the volume of the supercell and it contains only one doped atom. We consider the crystal structure representing MSG $P\bar{6}^{\prime}m^{\prime}2$ (No.187.211) [see Fig. 3(a)] as a prototype to study the effects of chemical doping in this category. Here, the doped atoms are placed on top of each other along $z$ and in alternate $xy$-planes. Therefore, in this case, it is possible to build the primitive unit cell by reducing the size of the supercell from $2\times 2\times 2$ to $2\times 2\times 1$. In this primitive unit cell, there exist four opposite spin pairs of magnetic atoms, and each pair creates a quasi-1D interaction block as shown by coloured cuboids in Fig.  8(a). Out of four such blocks, the nearest neighbor coordination of one block is not affected by the doping, and the corresponding chemical bonding nature remains unchanged (see the grey coloured cuboid), while the chemical bondings of other three blocks are affected by the doped atom (see the magenta, orange, and green blocks). The Hamiltonian for these three configurations is written as
@@
The basis order in the above Hamiltonian is $|\text{Mn}-d\rangle,\ |\text{Mn}^{\prime}-d\rangle,\ |\text{NM}_{\text{eff}}-p\rangle,\ |\text{NM}^{\prime}_{\text{eff}}-p\rangle$. Here, $|\text{NM}_{\text{eff}}-p\rangle$ is formed out of all Te atoms, while $|\text{NM}^{\prime}_{\text{eff}}-p\rangle$ includes one doped atom [see Fig. 8(a)]. The various hopping interaction terms in the Hamiltonian for the orange coloured interaction block are shown in Fig. 8(b) as an example. From the figure, we see that the coordination of the up-spin magnetic atoms is exactly the mirror image of that of the down-spin magnetic atoms. Therefore, the interactions among the magnetic and non-magnetic atoms obey the following relations: $w_{1}^{2}+w_{2}^{2}=v_{1}^{2}+v_{2}^{2}$ and $w_{3}^{2}+w_{4}^{2}=v_{3}^{2}+v_{4}^{2}$, and these relations lead to altermagnetic band structure [25], which is further validated by the band dispersions shown in Fig. 8(c).
Case II: Quasi-altermagnetism– As we discussed previously, chemical doping at any other site except the above-mentioned cases, resulted in two MSG families, namely $C2^{\prime}/m^{\prime}$ (No.12.62) and $P\bar{3}m^{\prime}1$ (No.164.89). In these cases, the ideal band dispersion defining altermagnetism disappears [see Fig. 3(b) and (d)]. To investigate the deviation from ideal altermagnetic band dispersion, we take the representative crystal structure of MSG $C2^{\prime}/m^{\prime}2$ (No.12.62) as a prototype. This system can be divided into four quasi-1D interaction blocks. Out of them, two are single doped and other two are pair doped. For the single doped interaction blocks, the characteristic AMSS is retained as discussed in the previous case.
@@
Here, the basis order is $|\text{Mn}_{1}-d\rangle,\ |\text{Mn}_{2}-d\rangle,\ |\text{Mn}_{3}-d\rangle,\ |\text{Mn}_{4}-d\rangle,\ |\text{NM}_{\text{eff}1}-p\rangle,\ |\text{NM}_{\text{eff}2}-p\rangle,\ |\text{NM}_{\text{eff}3}-p\rangle,\ |\text{NM}_{\text{eff}4}-p\rangle$. The triangular planes forming the effective orbitals $|\text{NM}_{\text{eff}1}-p\rangle,\ |\text{NM}_{\text{eff}2}-p\rangle$ have one doped atom. The other tringular planes forming the effective orbitals $|\text{NM}_{\text{eff}3}-p\rangle,\ |\text{NM}_{\text{eff}4}-p\rangle$ do not have any doped atoms. In Fig. 9(b), various hopping interactions in the Hamiltonian of Eq. 4 are schematically shown for the brown block as an example. The pair doped atoms introduce perturbations to the non-magnetic onsite energy ($\beta_{1}\neq\beta_{2}$), hopping strengths ($w_{1}+iw_{2}\neq w_{7}+iw_{8}$, $v_{1}+iv_{2}\neq w_{5}+iw_{6}$, etc.), and magnetic moments ($\Delta_{\text{Mn}}\neq\Delta^{\prime}_{\text{Mn}}$). Each perturbation independently lifts the antiferromagnetic sublattice band degeneracy. In the case of isovalent doping, such as with Se, the difference between $\beta_{1}$ and $\beta_{2}$ is much more prominent than the other two perturbations. In the case of hole-type (Sb) or electron-type (I) doping, each of the perturbations becomes prominent, leading to a larger deviation from the ideal altermagnetic band structure, which is substantiated through Fig. 9(d-f).

### v3

@@ III.4 Quasi-altermagnetism
Quasi-altermagnetism, as we define, refers to a magnetic phase that emerges when an altermagnet loses its rotational symmetries. This loss of symmetry can occur due to external stimuli such as strain, electric field, defects, and vacancies or chemical doping, as is the case in the present study. Like the altermagnets, quasi-altermagnets lack both translational symmetry and inversion symmetry. They do not exhibit Zeeman splitting as in the ferromagnets. Instead, they show an anisotropic momentum-dependent spin-splitting as in altermagnets. However, the splitting is not equal and opposite on either side of the nodal plane. In fact, the nodal plane itself is blurred and ill-defined in the case of quasi-altermagnets due to the fact that the sub-band degeneracy may get lifted on this plane. Furthermore, if the system is an insulator, the net magnetization remains zero, like the antiferromagnets and altermagnets. However, if it is metallic, it can have uncompensated magnetization. For example, MnTe1-xSex belongs to the SG $C2/m2$ (No. 12) and $P\bar{3}m1$ (No. 164) are insulators which exhibit quasi-altermagnetism with zero net magnetization. However, MnTe1-xSbx and MnTe1-xIx with the same MSG symmetry are quasi-altermagnetic metals, and these compounds have uncompensated magnetization. Fig. 5 illustrates the key features of quasi-altermagnetism and how they differ from ferromagnetism, antiferromagnetism, and altermagnetism.
While, altermagnetism requires perfect rotational symmetries, quasi-altermagnetism is a perturbative deviation effect. The constant energy plots of different pairs of bands in altermagnetic and quasi-altermagnetic materials, shown in Fig. 6, illustrate that quasi-altermagnetism originates from the breakdown in rotational symmetry of altermagnets, rather than the superposition of different types of altermagnetic order in varied proportion. In most of the practical systems, it is therefore highly likely that quasi-altermagnetism exists instead of altermagnetism.
In Appendix C, we have demonstrated the distinction among the various types of collinear magnetic ordering through schematic illustrations of their band structures.
@@ III.5 Anomalous Hall effect in MnTe with paired non-magnetic atom substitutions
The anomalous Hall effect (AHE) is a phenomenon that arises due to broken TRS and strong SOC. With the inclusion of SOC, altermagnets can exhibit AHC as the TRS is broken. However, the existence of AHE in altermagnets is not inherently guaranteed since it crucially depends on the orientation of the Néel vector and MSG symmetry associated with that orientation [17]. Hexagonal MnTe is one of the prominent altermagnetic materials in which AHE is both theoretically predicted and experimentally observed [17, 10]. In this section, we investigate AHE in MnTe and how it is modified by chemical doping.
In pristine MnTe, a finite AHC is obtained when Néel vector lies in the basal plane, orienting 30°  from crystal $a$-axis [17, 10]. The corresponding MSG in this configuration ($Cm^{\prime}c^{\prime}m$, No. 63.462) allows a hall vector in the $z$-direction, resulting a finite $\sigma_{xy}$ component. Interestingly, chemical doping does not suppress AHE in the in-plane Néel vector orientation. Although MSG of doped system differs from that of pristine, in all doped configuration symmetries allowing Hall vector present. A detailed symmetry based analysis about AHE when Néel vector orients in-plane is given in the Appendix D.
When the Néel vector is oriented in the out-of-plane direction, pristine MnTe does not exhibit AHE. In this configuration, AHC is symmetry-forbidden due to the presence of operations such as the three-fold rotations $C^{\pm}_{3z}$ and in-plane two-fold rotations $C_{210}$ and $C_{120}$ consistent with MSG $P6_{3}^{\prime}/m^{\prime}m^{\prime}c$ (No. 194.268). Similar to the pristine case, MnTe with either a single (non-magnetic) substitution or a pair (non-magnetic) substitution configuration with MSG $P\bar{6}^{\prime}m^{\prime}2$ (No. 187.211) does not exhibit AHE. It is because the symmetry operations that prohibit AHE remain preserved in these configurations. However, in the other pair doping configurations with reduced symmetry, AHE becomes symmetry allowed, leading to a finite AHE. For the pair doped configuration with MSG $P\bar{3}m^{\prime}1$ (No. 164.89), a Hall pseudo-vector along the $c$-axis becomes symmetry-allowed, as the orthogonal rotation axes combined with TRS ($\tau C_{2x}$, $\tau C_{2y}$) leave it invariant. In lower-symmetry structures with MSGs $Fm^{\prime}m^{\prime}2$ (No. 42.222) and $Pm^{\prime}m^{\prime}2$ (No. 25.60), the presence of the $C_{2z}$ symmetry operation enforces the vanishing of AHC components $\sigma_{yz}$ and $\sigma_{xz}$, while imposing no constraints on $\sigma_{xy}$. On the other hand, the lowest-symmetric and the most probable configuration with MSG $C2^{\prime}/m^{\prime}$ (No. 12.62) lacks any pure rotational symmetry and retains only the combined symmetry operation $\tau C_{2y}$, which allows non-zero components $\sigma_{yz}$ and $\sigma_{xy}$.
@@
In Fig. 7, we show the non-zero components of AHC and $\Omega$ for candidate materials in two distinct classes of pair non-magnetic atom doped MnTe systems – an altermagnet and a quasi-altermagnet – each incorporating Sb and I atoms as dopants. From the figure, we find that for MSG $Pm^{\prime}m^{\prime}2$ (No. 25.60) (both in Sb-doped [in (e)] and I-doped [in (g)] cases), the net contribution coming from $\Omega_{x}$ and $\Omega_{y}$ vanishes, as there exists always a negative counterpart for each of the positive Berry curvature. In contrast, $\Omega_{z}$ component has a net contribution over the entire Brillouin zone–contributing to non-zero $\sigma_{xy}$ as obtained from equation 1.
In a similar way, for MSG $C2^{\prime}/m^{\prime}$ (No. 12.62), in Fig. 7 (f) and (i), one can note that, only $\Omega_{y}$ component gets compensated, and other two ($\Omega_{x}$ and $\Omega_{z}$) remain uncompensated throughout the Brillouin zone. This gives rise to two non zero components of AHC – $\sigma_{yz}$ and $\sigma_{xy}$ – as depicted in Fig. 7 (b) and (d), respectively.
@@ III.6 Microscopic origin of altermagnetism in doped MnTe: A model Hamiltonian approach
where, i,ji,\ j and $\mu,\ \nu$ represent the site and orbital indices, respectively. In the Hamiltonian, the $d$-orbitals of the Mn atom and $p$-orbitals of the nonmagnetic atoms (Te, Se, Sb, I) are considered to construct the bases. Furthermore, $\sigma\in\{\uparrow,\downarrow\}$ and $\tau\in\{1,2\}$ indicate the spin and sublattice indices, respectively. In sublattice $\tau=1(2)$, the up (down) spin state dominates. In Eq. 2, the first term depicts the onsite energies of the individual orbitals. The second term represents antiferromagnetic Hund’s coupling, $\Delta$, of the spin-polarized orbitals (Mn $d$-orbitals in the present case). The (+)-(+) sign in this term corresponds to the $\downarrow(\uparrow)$ spin state of the electron. The sublattice index $\tau$ indicates that Mn ($\tau=1$) and Mn ($\tau=2$) sublattices have equal and opposite values of $\Delta/2$ for a given orbital. The third term represents the hopping of electrons, which are expressed using the SK relations. The detailed analysis for the chemical bonding origin of altermagnetism in NiAs prototype compounds can be found in Ref. [33]. Here, we primarily focus on the pair doping cases, as single doping is a subset of this. Furthermore, we intend to see the effect of doping in MnTe, which will affect only a part of the existing chemical bonding of the pristine crystal.
@@
If we examine the crystal structure of MnTe (for all NiAs prototypes) it can be presented as a series of chain like interaction blocks, as shown in Figs. 8(a) and 9(a). In the figures, each colored cuboid within the unit cell represents one such interaction block. If the electron-electron hopping interactions are restricted to the second nearest neighbor, these blocks are isolated from each other to behave like a “quasi-1D” chain. Therefore, the up- and down-spin blocks of the total Hamiltonian matrix can be written in terms of a set of block diagonal matrices with each block representing one of this quasi-1D chain. We note that by ignoring the interaction beyond the second neighbor, the altermagnetic characteristics are not affected  [33]. As the sole objective is to demonstrate how altermagnetic band dispersion gets affected in the presence of doping, for simplicity, we adapted a single orbital model (i.e., one $d$-orbital from the magnetic atom and one $p$-orbital from the non-magnetic atom). Earlier we have exclusively shown that it is possible to demonstrate altermagnetism using a single orbital picture. Within this framework, the basis of the Hamiltonian is now formed by the two $d$-orbitals, one each from the opposite spin sublattices, representing the magnetic site, and two effective hybridized orbitals representing the non-magnetic site. There are three nearest neighbor non-magnetic isoplanar atoms forming a triangle, one below a magnetic site and one above it [see right panel Fig. 8 (a)]. For each of the triangles, one effective hybridized orbital is formed out of the linear combination of three identical $p$-orbitals, coming from the non-magnetic atoms.
Figure 8: SK-TB analysis of altermagnetism in MnTe with pair doping. The primitive unit cell of the representative crystal of MSG $P\bar{6}^{\prime}m^{\prime}2$ (No. 187.211) [see Fig. 3(a)] is considered here. (a) Identification of different quasi-1D interaction blocks shown by color cuboids. From the electronic structure point of view, they are isolated from each other, if the hopping interactions are restricted up to the second nearest neighbors. In each quasi-1D block, a magnetic atom is capped by two iso-planar triangles, one above and one below, formed by the non-magnetic atoms. For the single-doped case, there exist four such quasi-1D interaction blocks, out of which one remains unaffected from doping (shown in gray) and the other three contain one doped atom Se/Sb/I. (b) Schematic illustration of the block diagonal tight-binding model Hamiltonian with each block representing one of the quasi-1D chains. (c) A visualization of the hopping interactions in the Hamiltonian matrix of one of the quasi-1D chains. (d) The band structure of configuration (c) is shown, where the altermagnetism is not affected by doping, with the characteristic AMSS retained.
@@
Figure 9: SK-TB analysis of quasi-altermagnetism in MnTe with pair doping. The quasi-1D interaction blocks are designed for the representative crystal of SG $C2/m$ (No. 12) as depicted in Fig. 3(b) to study the effect of doping. (a) There exist four different interaction blocks in the primitive unit cell in the framework of up to second neighbor interactions, which are shown by the color cuboids. Out of them, two interaction blocks contain only one doped atom (as shown by green and magenta colors in the upper panel), and the remaining two interaction blocks contain two doped atoms (as shown by brown and blue colors in the middle panel). While in the former case, the altermagnetic band structure is retained due to the presence of mirror symmetry, it vanishes, and a quasi-altermagnetic band structure is formed in the latter case because of the breaking of mirror symmetry. (b) Schematic illustration of the block diagonal tight-binding model Hamiltonian, with each block representing one of the quasi-1D chains. (c) The various hopping interactions are presented for a pair doped interaction block. Quasi-altermagnetic band structure for this configuration when there is a difference between the nonmagnetic onsite energies (d), hopping strengths (e), and local spin moments (f). In a real system, all three of them contribute together.
@@
Case I: Altermagnetism– Out of five different SG families of pair doped configurations, $P\bar{6}m2$ (No. 187), $Fmm2$ (No. 42), and $Pmm2$ (No. 25) come under the altermagnetism category, where a primitive unit cell exists with half the volume of the supercell and it contains only one doped atom. We consider the crystal structure representing SG $P\bar{6}m2$ (No. 187) [see Fig. 3(a)] as a prototype to study the effects of chemical doping in this category. Here, the doped atoms are placed on top of each other along $z$ and in alternate $xy$-planes. Therefore, in this case, it is possible to build the primitive unit cell by reducing the size of the supercell from $2\times 2\times 2$ to $2\times 2\times 1$. In this primitive unit cell, there exist four opposite spin pairs of magnetic atoms, and each pair creates a quasi-1D interaction block as shown by colored cuboids in Fig.  8(a). Out of four such blocks, the nearest neighbor coordination of one block is not affected by the doping, and the corresponding chemical bonding nature remains unchanged (see the gray colored cuboid), while the chemical bondings of other three blocks are affected by the doped atom (see the magenta, orange, and green blocks). The Hamiltonian for these three configurations is written as
@@
The basis order in the above Hamiltonian is $|\text{Mn}-d\rangle,\ |\text{Mn}^{\prime}-d\rangle,\ |\text{NM}_{\text{eff}}-p\rangle,\ |\text{NM}^{\prime}_{\text{eff}}-p\rangle$. Here, $|\text{NM}_{\text{eff}}-p\rangle$ is formed out of all Te atoms, while $|\text{NM}^{\prime}_{\text{eff}}-p\rangle$ includes one doped atom [see Fig. 8(a)]. The various hopping interaction terms in the Hamiltonian for the orange colored interaction block are shown in Fig. 8(b) as an example. The numerical values of the various interaction strengths used in Fig. 8(d) are listed in Table 2 of Appendix E. From the figure, we see that the coordination of the up-spin magnetic atoms is exactly the mirror image of that of the down-spin magnetic atoms. Therefore, the interactions among the magnetic and non-magnetic atoms obey the following relations: $w_{1}^{2}+w_{2}^{2}=v_{1}^{2}+v_{2}^{2}$ and $w_{3}^{2}+w_{4}^{2}=v_{3}^{2}+v_{4}^{2}$, and these relations lead to altermagnetic band structure [33], which is further validated by the band dispersions shown in Fig. 8(c).
Case II: Quasi-altermagnetism– As we discussed previously, chemical doping at any other site except the above-mentioned cases, resulted in two SG families, namely $C2/m$ (No. 12) and $P\bar{3}m1$ (No. 164). In these cases, the ideal band dispersion defining altermagnetism disappears [see Fig. 3(b) and (d)]. To investigate the deviation from ideal altermagnetic band dispersion, we take the representative crystal structure of SG $C2/m$ (No. 12) as a prototype. This system can be divided into four quasi-1D interaction blocks. Out of them, two are single doped and other two are pair doped. For the single doped interaction blocks, the characteristic AMSS is retained as discussed in the previous case.
@@
Here, the basis order is $|\text{Mn}_{1}-d\rangle,\ |\text{Mn}_{2}-d\rangle,\ |\text{Mn}_{3}-d\rangle,\ |\text{Mn}_{4}-d\rangle,\ |\text{NM}_{\text{eff}1}-p\rangle,\ |\text{NM}_{\text{eff}2}-p\rangle,\ |\text{NM}_{\text{eff}3}-p\rangle,\ |\text{NM}_{\text{eff}4}-p\rangle$. The triangular planes forming the effective orbitals $|\text{NM}_{\text{eff}1}-p\rangle,\ |\text{NM}_{\text{eff}2}-p\rangle$ have one doped atom. The other triangular planes forming the effective orbitals $|\text{NM}_{\text{eff}3}-p\rangle,\ |\text{NM}_{\text{eff}4}-p\rangle$ do not have any doped atoms. In Fig. 9(b), various hopping interactions in the Hamiltonian of Eq. 4 are schematically shown for the brown block as an example (see Table 2 of Appendix E for the numerical values of interaction strengths). The pair doped atoms introduce perturbations to the non-magnetic onsite energy ($\beta_{1}\neq\beta_{2}$), hopping strengths ($w_{1}+iw_{2}\neq w_{7}+iw_{8}$, $v_{1}+iv_{2}\neq w_{5}+iw_{6}$, etc.), and magnetic moments ($\Delta_{\text{Mn}}\neq\Delta^{\prime}_{\text{Mn}}$). Each perturbation independently lifts the antiferromagnetic sublattice band degeneracy. In the case of isovalent doping, such as with Se, the difference between $\beta_{1}$ and $\beta_{2}$ is much more prominent than the other two perturbations. In the case of hole-type (Sb) or electron-type (I) doping, each of the perturbations becomes prominent, leading to a larger deviation from the ideal altermagnetic band structure, which is substantiated through Fig. 9(d-f).